OpenStudy (anonymous):

can someone show me step by step how to do this? If $360 is invested at an interest rate of 4% per year and is compounded quarterly, how much will the investment be worth in 18 years?

OpenStudy (anonymous):

Do you know the compound interest equation?

OpenStudy (anonymous):

your ormuila is A= P(1+r/n)^nt where in this case n= 4 , r=4%, P=$360 , t= 18 yeras

OpenStudy (anonymous):

i hope this will help you

OpenStudy (anonymous):

yes I do , but I try it and it gave me a weird answer

OpenStudy (anonymous):

I got 172.657?

OpenStudy (anonymous):

is there rounding that I have to do?

OpenStudy (anonymous):

$736.95 is the ans

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

736.96 is one of my choices.

OpenStudy (anonymous):

ya thats the ans

OpenStudy (anonymous):

but see I got 172.65? and I wanted to know where I messed up , but nm thanks

OpenStudy (anonymous):

see the calculation part A= 360(1+.04/4)^4x18 =360(1+.01)^72 =360(1.01)^72 =360x2.0471=736.955

OpenStudy (anonymous):

did you get it now